OpenStudy (anonymous):

HELP ME PLEASE Solve for y: y over negative 2+ 5 = −4

OpenStudy (ashleyisakitty):

\[\frac{y }{ 2+5}=-4\]

OpenStudy (ashleyisakitty):

Is this your equation?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

y over -2 + 5

OpenStudy (ashleyisakitty):

\[\frac{ y }{ -2+5 }=-4\]

OpenStudy (ashleyisakitty):

Is this correct? Your title is broken so I cannot tell.

OpenStudy (anonymous):